# Vulnerability Webs:

### v2

@@ 5 Contagious Vulnerabilities
A growing literature explores contagion on networks, including the statistical literature on infectious diseases (e.g., the contagion of HIV and coronaviruses). Such models (see, e.g., Schweinberger et al., 2022; Groendyke et al., 2012; Britton and O’Neill, 2002) usually first use a network formation model to generate a network of contacts among agents and then study how an infectious disease following a stochastic process—like the susceptible-infected-recovered model (Kermack and McKendrick, 1972; Hethcote, 2000)—spreads within this network.
We adapt these ideas to the contagion of vulnerabilities in software dependency graphs. The prevalence of vulnerabilities in software motivated the creation of the Common Vulnerabilities and Exposure (CVE) program and the National Vulnerability Database (NVD) by the National Institute of Standards and Technology, which is an agency of the United States Department of Commerce (NIST, 2024). A software package depending on a vulnerable package is potentially vulnerable itself. To see how, consider, for example, the Equifax data breach (Federal Trade Commission, 2024) in 2017, caused by a vulnerability in the popular open source web application framework Apache Struts 2.262626Apache Struts 2 did not properly validate the Content-Type header of incoming http requests when processing file uploads, which allowed attackers to create http requests that include malicious code. Due to the lack of content validation, this allowed attackers to execute arbitrary code on the web server, ultimately granting them full control of the server. For details see the official NIST announcement on CVE-2017-5638 (https://nvd.nist.gov/vuln/detail/CVE-2017-5638). Equifax did not itself develop or maintain Apache Struts 2, but instead used it as part of its own code base. The vulnerability allowed attackers to execute malicious code providing access to the server hosting the web application developed by Equifax using Apache Struts 2 and costing the company more than \$1.5 Billion to date. [^27] See https://www.bankinfosecurity.com/equifaxs-data-breach-costs-hit-14-billion-a-12473. Article accessed 27 January 2024.
Package managers are part of a large suite of tools that help systems administrators to keep their systems up to date and patch any known vulnerabilities. It takes time for vulnerable systems to be updated (Edgescan, 2022), [^30] It takes organizations between 146 and 292 days to patch cyber security vulnerabilities (Source: missing
@@ 6 Conclusion
Appendix A Computational details for estimation
Estimation of our model is challenging because of the normalizing constants $c_{kk}$ in the likelihood and the discrete mixture model for the block assignments. We bypass these issues by extending recently developed approximate two-step estimation methods to directed networks, estimating the structural parameters in two steps.
Formally, the full likelihood of our model can be written as follows
$\mathcal{L}(\bm{g},\bm{x};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta})=\sum_{\bm{z}\in\mathcal{Z}}L\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta}\right)=\sum_{\bm{z}\in\mathcal{Z}}p_{\bm{\eta}}\left(\bm{Z}=\bm{z}\right)\,\pi(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma).$ (18)
where each package’s type is i.i.d. multinomial, so the distribution $p_{\bm{\eta}}(\bm{z})$ is
[…]
Since maximizing (18) involves a sum over all possible type allocations of each node as well as a nested normalizing constant in $\pi(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma)$, evaluating the function is infeasible. To ameliorate this issue, we divide estimation into two parts. First, we estimate the type of each node and, second, we estimate the parameters $\bm{\alpha},\bm{\beta},$ and $\gamma$ conditional on the type of each node. This two-step algorithm exploits the fact that our model corresponds to a stochastic blockmodel with covariates when $\gamma=0$, i.e. when there are no link externalities.
A.1 STEP 1: Approximate estimation of unobserved block structure
A.1.1 Variational EM algorithm with MM updates
To recover the types and the block structure of the network, we approximate the model using a stochastic blockmodel. Define
$L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right):=p_{\eta}(\bm{z})\pi(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma=0)$ (19)
as the likelihood of a stochastic blockmodel, which means that
  1. 1.
    Each node belongs to one of $K$ blocks/types;
  2. 2.
    Each link is conditionally independent, given the block structure
Then under some conditions – namely, that the network is large, each block/type is not too large with respect to the network, and the employed statistics are in some sense well-behaved (Babkin et al., 2020; Schweinberger, 2020; Schweinberger and Stewart, 2020) – we have
$L\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta}\right)\approx L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right).$ (20)
Denoting by $q_{\xi}(\bm{z})$ a the auxiliary distribution characterized by parameter $\bm{\Xi}$ approximating the distribution $p_{\eta}(\bm{z})$, the log-likelihood of our model stated in (18) can be lower bounded as follows:
$\ell(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma,\bm{\eta})$ ::=:= $\log\sum_{\bm{z}\in\mathcal{Z}}L\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta}\right)$ (21)
$\approx$ $\log\sum_{\bm{z}\in\mathcal{Z}}L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)$
$=$ $\log\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\frac{L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)}{q_{\xi}(\bm{z})}$
$\geq$ $\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\log\left(\frac{L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)}{q_{\xi}(\bm{z})}\right)$
$=$ $\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\log L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)-\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\log q_{\xi}(\bm{z})$
$=$ $\mathbb{E}_{q}\log L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)+\mathbb{H}(q)$
$\eqqcolon$ $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi}).$
@@ A.1.1 Variational EM algorithm with MM updates
$\mathbb{H}(q)=-\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\log q_{\xi}(\bm{z})$
is the entropy of auxiliary distribution $q_{\xi}(\bm{z})$. In the third row, we multiply and divide by $q_{\xi}(\bm{z})$ and apply Jensen’s inequality in the consecutive line.
The best lower bound is obtained by choosing $q_{\xi}(z)$ from the set of distributions $\mathcal{Q}$ that solves the following variational problem
$\ell(\bm{\alpha},\bm{\beta},\bm{\eta})=\sup_{\bm{\Xi}\,\in\,[0,1]^{N\times K}}\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})$
However, this variational problem is usually intractable unless we impose more structure on the problem. In practice, researchers restrict the set $\mathcal{Q}$ to a smaller set of tractable distributions (Wainwright and Jordan, 2008; Mele and Zhu, 2023). In the case of the stochastic blockmodel, it is useful and intuitive to restrict $\mathcal{Q}$ to the set of multinomial distributions
[…]
where we collect the variational parameters in the matrix $\bm{\Xi}=(\xi_{ik})\in[0,1]^{N\times K}$. This leads to the following tractable lower bound of $\ell$
$\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})$$\equiv$ $\sum_{\bm{z}\in\mathcal{Z}}q_{\bm{\Xi}}(\bm{z})\log\left(\frac{L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)}{q_{\bm{\Xi}}(\bm{z})}\right)$
$=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\xi_{ik}\xi_{jl}\log\,\pi_{ij,kl}(g_{ij},\bm{x})$
$+$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\xi_{ik}\left(\log\eta_{k}-\log\xi_{ik}\right),$
where the function $\pi_{ij,kl}$ is the conditional probability of a link between $i$ and $j$ of types $k$ and $l$, respectively,
$\log\,\pi_{ij,kl}(g_{ij},\bm{x})$$\coloneqq$ $g_{ij}\log\left(\frac{\exp\left(u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right)}{1+\exp\left(u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right)}\right)$
$+$ $(1-g_{ij})\log\left(\frac{1}{1+\exp\left(u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right)}\right)$

### v3

@@ 5 Contagious Vulnerabilities
A growing literature explores contagion on networks, including the statistical literature on infectious diseases (e.g., the contagion of HIV and coronaviruses). Such models (see, e.g., Schweinberger et al., 2022; Groendyke et al., 2012; Britton and O’Neill, 2002) usually first use a network formation model to generate a network of contacts among agents and then study how an infectious disease following a stochastic process spreads within this network.
We adapt these ideas to the contagion of vulnerabilities in software dependency graphs. The prevalence of vulnerabilities in software motivated the creation of the Common Vulnerabilities and Exposure (CVE) program and the National Vulnerability Database (NVD) by the National Institute of Standards and Technology, which is an agency of the United States Department of Commerce (NIST, 2024). A software package depending on a vulnerable package is potentially vulnerable itself. To see how, consider, for example, the Equifax data breach (Federal Trade Commission, 2024) in 2017, caused by a vulnerability in the popular open source web application framework Apache Struts 2.252525Apache Struts 2 did not properly validate the Content-Type header of incoming http requests when processing file uploads, which allowed attackers to create http requests that include malicious code. Due to the lack of content validation, this allowed attackers to execute arbitrary code on the web server, ultimately granting them full control of the server. For details see the official NIST announcement on CVE-2017-5638 (https://nvd.nist.gov/vuln/detail/CVE-2017-5638). Equifax did not itself develop or maintain Apache Struts 2, but instead used it as part of its own code base. The vulnerability allowed attackers to execute malicious code providing access to the server hosting the web application developed by Equifax using Apache Struts 2 and costing the company more than \$1.5 Billion to date. [^26] See https://www.bankinfosecurity.com/equifaxs-data-breach-costs-hit-14-billion-a-12473. Article accessed 27 January 2024.
Package managers are part of a large suite of tools that help system administrators keep their systems up to date and patch any known vulnerabilities. It takes time for vulnerable systems to be updated (Edgescan, 2022), [^28] It takes organizations between 146 and 292 days to patch cyber security vulnerabilities (Source: missing
@@ 6 Conclusion
Appendix A Exploiting sparse matrix operations for updating variational parameters
Next, we detail how the updates of $\mathbf{\xi}$ in Step 1 can be carried out in a scalable manner based on sparse matrix operations. Equation (17) can be written as a quadratic programming problem in $\mathbf{\xi}$ with the block constraints that $\sum_{k=1}^{K}\xi_{ik}=1$ needs to hold for all $i=1,\ldots,N$. We rearrange the first line of (17) as follows:
$\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\,\frac{\xi_{jl}^{(t)}}{2\xi_{ik}^{(t)}}+\xi_{jl}^{2}\,\frac{\xi_{ik}^{(t)}}{2\xi_{jl}^{(t)}}\right)\log\,\pi_{kl}^{(t)}(g_{ij},\bm{x}_{ij})$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{ik}^{2}\,\frac{\xi_{jl}^{(t)}}{2\xi_{ik}^{(t)}}\left(\log\,\pi_{kl}^{(t)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(t)}(g_{ji},\bm{x}_{ji})\right)$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}\frac{\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(t)})}{2\xi_{ik}^{(t)}}\,\xi_{ik}^{2}$
with
$\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)~{}\coloneqq~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(t)}\left(\log\,\pi_{kl}^{(t)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(t)}(g_{ji},\bm{x}_{ji})\right).$ (19)
We make the dependence of Equation (19) on $\bm{g}$ and $\bm{x}$ specific to ease the notation at later steps of this derivation. This yields for Equation (17)
$M\left(\bm{\Xi};\bm{\theta},\bm{\eta}^{(t)},\bm{\Xi}^{(t)}\right)$ $=\sum_{i=1}^{N}\sum_{k=1}^{K}\left(\frac{\Omega_{ik}^{(t)}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)}{2\,\xi_{ik}^{(t)}}-\frac{1}{\xi_{ik}^{(t)}}\right)\xi_{ik}^{2}+\left(\log\eta_{k}^{(t)}-\log\xi_{ik}^{(t)}+1\right)\xi_{ik}$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)\,\xi_{ik}^{2}+B_{ik}\left(\bm{\eta}^{(t)},\bm{\Xi}^{(t)}\right)\,\xi_{ik}$
@@ Appendix A Exploiting sparse matrix operations for updating variational parameters
$A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)\coloneqq\frac{\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)}{2\,\xi_{ik}^{(t)}}-\frac{1}{\xi_{ik}^{(t)}}$
is the quadratic term and
$B_{ik}\left(\bm{\eta}^{(t)},\bm{\Xi}^{(t)}\right)\coloneqq\log\eta_{k}^{(t)}-\log\xi_{ik}^{(t)}+1$
the linear term of the quadratic problem.
To update the estimate of $\bm{\Xi}$, we need to evaluate $A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)$ and $B_{ik}\left(\bm{\eta}^{(t)},\bm{\Xi}^{(t)}\right)$ for $i=1,\ldots,N$ and $k=1,\ldots,K$, which, when done naively, is of complexity $O(N^{2}\,K^{2})$. Computing $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)$ for $A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)$ is the computational bottleneck driving these complexities. It is thus prohibitively difficult for a large number of population members and communities in the network. Note that in our application, this problem is exasperated as we have more than 35,000 nodes and our theoretical result on the adequacy of the estimation procedure assumes that $K$ grows as a function of $N$. To avoid this issue, we show how $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)$ can be evaluated through a series of sparse matrix multiplications. Thereby, our implementation is fast and light on memory requirements. The underlying idea is that $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)$ collapses to a simple form if $\mathbf{g}$ is an empty network with all pairwise covariates set to zero, i.e., $g_{ij}=0$ and $x_{q}=0\,\forall\,i,j=1,\ldots,N$ and $q=1,\ldots,p$. Given that, in reality, these are seldom the observed values, we, consecutively, go through all connections where either $g_{ij}=1$ or $X_{ij,1}=x_{1},\ldots,X_{ij,p}=x_{p}$ for $\sum_{q=1}^{p}\,x_{q}\neq 0$ holds and correct for the resulting error. In formulae, this implies the following:
$\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(t)})~{}$ $\coloneqq~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(t)}\left(\log\,\pi_{kl}^{(t)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(t)}(g_{ji},\bm{x}_{ji})\right)$
$=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(t)}\left(\log\,\pi_{kl}^{(t)}(0,\bm{0})+\log\,\pi_{lk}^{(t)}(0,\bm{0})\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{kl}^{(t)}(1,\bm{x}_{ij})}{\pi_{kl}^{(t)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{lk}^{(t)}(1,\bm{x}_{ji})}{\pi_{lk}^{(t)}(0,\bm{0})}\right)\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{kl}^{(t)}(0,\bm{x}_{ij})}{\pi_{kl}^{(t)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{lk}^{(t)}(0,\bm{x}_{ji})}{\pi_{lk}^{(t)}(0,\bm{0})}\right)\right)$
$=\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(t)}\right)+\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right),$ (20)
with
$\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)$ $=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{kl}^{(t)}(1,\bm{x}_{ij})}{\pi_{kl}^{(t)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{lk}^{(t)}(1,\bm{x}_{ji})}{\pi_{lk}^{(t)}(0,\bm{0})}\right)\right)$ (21)
$+\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{kl}^{(t)}(0,\bm{x}_{ij})}{\pi_{kl}^{(t)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{lk}^{(t)}(0,\bm{x}_{ji})}{\pi_{lk}^{(t)}(0,\bm{0})}\right)\right)$
being the error in $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)$ arising from assuming an empty network with all categorical covariates set to zero. Next, we show how both terms, $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(t)}\right)$ and $\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)$, can be computed through matrix operations, completing our scalable algorithmic approach.
We, first, assume an empty network where all categorical covariates set to zero, $g_{ij}=x_{q}=0$ for all $i\neq j$ and $q=1,...,p$ and compute $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(t)}\right)$ via matrix multiplications. Observe that
$\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(t)}\right)$ $=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(t)}\left(\log\,\pi_{kl}^{(t)}(0,\bm{0})+\log\,\pi_{lk}^{(t)}(0,\bm{0})\right)$
$=\sum_{l=1}^{K}\sum_{j\neq i}^{N}\xi_{jl}^{(t)}\left(\log\,\pi_{kl}^{(t)}(0,\bm{0})+\log\,\pi_{lk}^{(t)}(0,\bm{0})\right)$
$=\sum_{l=1}^{K}\left(\sum_{j=1}^{N}\xi_{jl}^{(t)}-\xi_{il}^{(t)}\right)\left(\log\,\pi_{kl}^{(t)}(0,\bm{0})+\log\,\pi_{lk}^{(t)}(0,\bm{0})\right)$
$=\sum_{l=1}^{K}\left(\tau^{(t)}_{l}-\xi_{il}^{(t)}\right)\left(\log\,\pi_{kl}^{(t)}(0,\bm{0})+\log\,\pi_{lk}^{(t)}(0,\bm{0})\right)$
holds, where
$\tau^{(t)}_{l}~{}\coloneqq~{}\sum_{j=1}^{N}\xi_{jl}^{(t)}.$
With
$\bm{A_{0}}\left(\bm{\Xi}^{(t)}\right)~{}\coloneqq~{}\begin{bmatrix}\tau^{(t)}_{1}-\xi_{11}^{(t)}&\tau^{(t)}_{2}-\xi_{12}^{(t)}&\dots&\tau^{(t)}_{K}-\xi_{1K}^{(t)}\\ \tau^{(t)}_{1}-\xi_{21}^{(t)}&\tau^{(t)}_{2}-\xi_{22}^{(t)}&\dots&\tau^{(t)}_{K}-\xi_{2K}^{(t)}\\ \vdots&\vdots&\ddots&\vdots\\ \tau^{(t)}(1)-\xi_{n1}^{(t)}&\tau^{(t)}_{2}-\xi_{n2}^{(t)}&\dots&\tau^{(t)}_{K}-\xi_{nK}^{(t)}\end{bmatrix}$